OpenStudy (anonymous):

Simplify (-5m^2q)^2(-3m^3q)^3

OpenStudy (anonymous):

Is that (-5m to the power of 2q) or (-5m to the power of 2) times q? If it's (-5m to the power of 2) times q then here's your working: (-5m^2q)^2(-3m^3q)^3 (-5m^4q^3)(-3m^6q^4) -5m^4*-3m^6=15m^10 q^3*q^4=2q^7 (-5m^2q)^2(-3m^3q)^3=15m^10*2q^7

OpenStudy (anonymous):

But the options are A.675m13q5 B.-675m5q2 C.-675m13q5 D.675m5q2